Steel production from a mill generates a negative externality because of the environmental damage linked to air and water pollution.  Suppose the market demand and supply curves are given by:

Demand (MB):  P = 400 - 3Qd

Supply (MC):  P = 200 + Qs

Q is tons of steel and P is price per ton of steel.

  

Note in this form, the demand and supply curve are solved for P -- you can see directly the lines on our supply and demand graph.  So, the intercept of demand curve (MB) is $400 and intercept of supply (MC) curve is 200!

Important: Each ton of steel creates a marginal external cost (MEC) of $80.

(a) Calculate the competitive market equilibrium price and quantity.  This would be solving for Qmarket, as in book or videos.

(b) Calculate the optimal quantity of steel production when you utilize the marginal social cost (MSC).  Remember the MSC=MC+MEC.  Hint: add MEC value to intercept of the supply curve to obtain the MSC curve.

(c) Calculate the total external cost (TEC) at both the competitive market output level and the optimal output level. TEC = MEC*Q for each case.   Provide a brief explanation of the results.  Is it optimal to have some pollution in this example?

(d) On a graph, clearly label the deadweight loss that occurs when production is at the market equilibrium you calculated in part (a).  Explain why this is not the level of output that maximizes social surplus.
Answer #1

A) Competitive equilibrium at where market demand EQUALS to market supply.

400-3Q=200+Q

200=4Q

Q=200/4=50

P=400-3*50=250

B) optimal QUANTITY is where price Equal to Marginal social cost.

MSC=MC+MEC=200+q+80=280+q

Optimal QUANTITY,

280+q=400-3q

-120=-4q

Q=120/4=30

C) because Marginal external cost is fixed and constant ,so,

total external cost=MEC*Q

At competitive equilibrium,

Total external cost=80*50=400

At optimal QUANTITY,

Total external cost=80*30=240

Yes , producing a good gives CONSUMERs utility and also generate pollution which generate external cost.so till external cost plus Production cost is equal to marginal utility , having pollution is optimal. Because the net benefit ( benefit- costl is maximized.

D)MSC ME optimet Quantity deadweight loss 2008 competitive equilibonim pie 2007 3050 Onentta

The demand curve shows marginal benefit while MSC shows marginal cost of good.

At optimal QUANTITY, marginal social cost of good is equal to marginal benefit of the good ,so higher or lower Production will decrease net benefit.

At competitive equilibrium , demand curve is below,MSC curve ,it means marginal benefit is lower than marginal social cost.,so net benefit is not Maximizing.

The further will we go from Optimal QUANTITY,the larger the gap between marginal cost and marginal benefit will be and the lower will net benefit.